List the members of these sets.

a) {x | x is a real number such that x 2 = 1}

b) {x | x is a positive integer less than 12}


1
Expert's answer
2022-03-19T02:39:47-0400

Solution (a)


Given that


"Set\ A=\left \{ x | x \text \ is\ a\ real\ number\ such\ that\ x^2 = 1 \right \}"


solving


"x^2=1"


"x=\pm1"


Hence the given set in the tabular form will be


"Set\ A=\left \{ 1, -1\right \}"


Solution (b)


Given that


"Set\ B=\left \{ x | x\ is\ a\ positive\ integer\ less\ than\ 12 \right \}"


Clearly we know the positive integers less than 12 are "1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11"


Therefore, the given set in the tabular form will be


"Set\ B=\left \{ 1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11\right \}"